How much time will be required for a sample of H-3 to lose 75% of its radioactivity? The half-life of tritium is 12.26 years.?





1
Expert's answer
2021-12-06T15:47:04-0500

Solution:

Tritium = H-3

The half-life (t1/2) of tritium is 12.26 years.


From law of radioactive decay:

t1/2 = ln(2) / λ

where λ = decay constant

λ = ln(2) / t1/2

λ = (0.693) / (12.26 years) = 0.0565 years-1


Radioactive decay follows the kinetics of a first order reaction:

ln[H-3]t = -λt + ln[H-3]o


If the sample lose 75% (0.75), then 25% (0.25) remains.

[H-3]t = 0.25

[H-3]o = 1.00

λ = 0.0565 years-1

Therefore,

ln(0.25) = - (0.0565 years-1) × t + ln(1.00)

-1.3863 = - (0.0565 years-1) × t

t = (-1.3863) / (-0.0565) = 24.54 years

t = 24.54 years


Answer: 24.54 years will be required for a sample of H-3 to lose 75% of its radioactivity.